Please help me. I am having problem solving these question.
1. 3 + 1/x-2 = 2x-3/x-2
2. 4/a^2-1 + 1/2a+2
Thank you

Answer by stanbon(48544)

1. 3 + 1/(x-2) = (2x-3)/(x-2)
--------
Comment: When you have an equation with denominators multiply
thru by the least common multiple of the denominators to get
rid of the denominators.
------------------------
Your Problem:
Multiply thru by (x-2) to get:
3(x-2) + 1 = (2x-3)
3x-6+1 = 2x-3
3x-5 = 2x-3
x = 2
===========

2. 4/(a^2-1) + 1/(2a+2)
----------
Comment: When you add or subtract fractions you need
to rewrite each fraction over the least common divisor
of the denominators.
-------------
Your Problem:
Factor where you can to get:
4/[(a-1)(a+1)] + 1/[2(a+1)]
The lcd = 2(a+1)(a-1)
Rewrite each fraction:
4*2/lcd + 1*(a-1)/lcd
Combine the numerators over the lcd.
[8+a-1]/lcd
(a+7)/[2(a-1)(a+1)]
==========
Cheers,
Stan H.